: Propane burns according the following reaction
C3H8 + 5O2 3CO2 + 4H2O
Calculate the following
1)How many moles of C3H8 are needed to react with 0.36 mole of O2
2) How many moles of CO2 can form from 0.36 mole of O2
3) How many moles of O2 are required for the reaction of 0.72 mole of C3H8
4) How many moles of H2O can form 0.36 moles O2
1
Expert's answer
2020-12-16T07:52:07-0500

1) Find the amount of substance of C3H8

.n(C3H8)= 1/5 n(O2)= 0.36/5= 0.072 mol

2) n (CO2)=3/5 n(O2)= 3/5 *0.36=0.216 mol

3)n (O2)=5 n(C3H8)=0.72 *5=3.6 mol

4) n(H2O)=4/5 n(O2)=4/5 *0.36=0.288 mol